# Public-private funding models in open source software development: A case study on scikit-learn

### v3

@@ 3.2 Case presentation
The landscape of OSS in the field of ML/AI has evolved significantly in the last two decades. In 2007, a group of leading scientists lamented the dearth of OSS for ML and launched the Journal of Machine Learning Open Source Software to support OSS development in this scientific field [77]. Nowadays, OSS is considered indispensable to ML/AI R&D [78, 79] and, congruent with industry dominance in AI R&D [10], a significant number of OSS libraries for ML/AI have been initiated, hosted, and maintained by industry actors [80]. However, a noteworthy exception to this trend is scikit-learn, an OSS project that was born out of Inria, a public science body in France, and that is supported by a global community of volunteers. The following section provides an overview of the scikit-learn project.
scikit-learn is a Python library that implements algorithms for classification, regression, and clustering, as well as related tools for data preprocessing, model fitting, model evaluation, and visualisations. Initially called scikits.learn, the project was started by David Courapeau as a Google Summer of Code project in 2007. After a dormant period, it was relaunched as scikit-learn by Fabian Pedregosa, Gaël Varoquaux, Alexandre Gramfort, and Vincent Michel with the backing of Inria, with the first public release in February 2010. Nowadays, it is one of the most impactful and popular projects in the ML/AI OSS landscape. It is described as “the Swiss army knife of ML” due to its widespread use in research and industry [9]. By January 2022, it boasted a community of 2,250 contributors, 46,177 GitHub stars, and the most downloads (621,408,470) out of all Python libraries in this landscape between January 2017 and December 2021. By November 2023, the project boasted 2,752 contributors and 61,085 GitHub stars [81].
During the interviews, the maintainers explained that the sustainability of scikit-learn was not inevitable. There were many uncertainties throughout the project’s history, from the under-development of the ML OSS ecosystem in the early 2010s to market pressure in the later 2010s. One factor, amongst others, that contributed to the project’s sustainability was its funding model. The project has benefitted from several sources of funding [82]. Inria is the longest-standing funder, which has paid for the salaries of maintainers, office space, and events, amongst others, since 2010. In 2018, the scikit-learn consortium was established under Inria Foundation, comprising companies that pay an annual membership fee, in order to secure employment for maintainers, sustain the quality of software, and develop new features [83]. Since its founding, the consortium has included Dataiku, Microsoft, Nvidia, Intel, AXA, Boston Consulting Group, BNP Paribas Cardif, Hugging Face, Nvidia, Fujitsu, and Chanel. Other companies, such as the OSS consultancy Quansight Labs, have indirectly funded the project by sponsoring scikit-learn maintainers. The project also benefits from donations from community members via NUMFOCUS and student projects have been sponsored by the Google Summer of Code programme. In November 2021, the French government announced a €32 million grant in its AI strategy for the expansion of scikit-learn and the development of open source tools for data science [11]. The mixed funding of scikit-learn makes it a suitable case for a targeted analysis of the aforementioned research questions.
@@ 3.3.2 Qualitative data analysis
Throughout the 17 months of data collection, an iterative, integrated approach was employed to identify common themes in the interview data, field notes, and secondary documents [85]. Given the exploratory aims of this study, the integrated approach combined the application of grounded theory analysis [86] and a deductive, organising framework informed by prior literature about the different interests and involvement models of individuals, companies, and governments in the development of OSS [40, 58, 73, 24]. Thematic codes were continuously developed as data was collected, and in turn translated into themes until the point of saturation [87]. The resulting themes were member-checked to ensure their accuracy and resonance with research participants [88]. Throughout the research process, a social identity map was maintained by the author to exercise reflexivity regarding the potential influence of personal biases stemming from the author’s views and relationships built ith the scikit-learn team during the 17 months of data collection [89].
@@ 4 Results
The results are organised into three parts. The first part discusses the interests and stakes of the scikit-learn maintainers, while the second and third parts respectively report on the interests and involvement of the public and private funders. This structure sequentially responds to RQ1 concerning how the interests of the project’s funders align or conflict with those of the maintainers, and RQ2 concerning the relative merits and drawbacks of the different types of funding. With regards to RQ3, practical lessons from this analysis are presented in the subsequent Discussion section.
4.1 Stewards of the community: The incentives and stakes of the maintainers
The maintainers of scikit-learn are a heterogeneous group, comprising employees at Inria based at the Saclay science park on the outskirts of Paris and active community members, who, at the time of this research, were employed by companies, such as Nvidia, Quansight Labs, Microsoft, and Hugging Face. Several maintainers also held academic positions at universities and research labs. While some members of the core team worked on scikit-learn full-time, others balanced their roles as maintainers with their jobs. One maintainer explained that the maintainers, who are not employed by Inria, have varying levels of capacity to contribute because they have to balance competing demands on their time. Furthermore, the maintainers have had different journeys to scikit-learn. While some worked their way up to the role of maintainer as employees at Inria, others were invited based on their active involvement as community members. For example, Adrin Jalali explained that he began contributing in 2018 because he used scikit-learn as a data science consultant and got increasingly involved because “I got good feedback and there was a good community, so I started spending hours and hours and weekends on the project, and about six months later, I woke up to an email asking me if I wanted to be a core contributor.”
Maintainers shared the conviction that scikit-learn would not be what it is today without its community, which one maintainer argued was at the heart of the project’s DNA. During an on-site visit to the scikit-learn office, François Goupil, the community manager, showed me a dashboard of activity in the project, pointing to world maps of contributors. While the majority of the maintainers are based in Paris, France, the community contains contributors from all over the world (see Table 1). The maintainers lay great importance on this diversity and make efforts to cultivate the project’s identity as the collective effort of a global community. The dashboard also showed the involvement of various organisations, from research institutes and universities to startups and industry giants (see Table 2). Upon observing that the most active organisations are Chinese and US American Big Tech companies, I asked François if knowing how active these companies are had influenced his opinion about whether they contribute enough to the project. He responded frankly: while such companies contribute useful issues and pull requests, at the end of the day their contributions create more work for the already overloaded maintainers. It would be more helpful if they sponsored a developer or joined the consortium in order to provide more capacity to process the extra workload created by their contributions.
@@ 4.1 Stewards of the community: The incentives and stakes of the maintainers
The maintenance of scikit-learn is a multifaceted job, spanning expertise in software engineering and statistical theory as well as social tasks like contributor mentorship and community animation, amongst others. François explained, “Of course, software development itself is the major part of the workload—going through issues, pull requests, and responding to community feedback—but lots of other work goes into the project. Community animation is a huge part. After all, it’s a community project.” He elaborated that in some ways the maintainers act like “editors” since they do not rush to implement issues or pull requests; rather, they carefully assess the needs and significance of issues and pull requests, liaise with maintainers of key dependencies, attend conferences and workshops, and read papers to stay up to date with research trends. “We don’t want to make easy, small fixes that crash in a few months,” he explained. Furthermore, sponsored developers from companies make focused contributions, such as the development of new features, which are collectively agreed upon by the maintainers and the sponsoring company. For example, at the time of this research, Nvidia sponsord a full-time maintainer who was building a plug-in to enable alternative computational backends for scikit-learn to enable compatibility with Nvidia’s AI hardware accelerators.
The demands of the job begs the question of incentives. Some maintainers explained that the work resonated with their values. The first time I met Gaël Varoquaux, he described scikit-learn as a political project that aims to provide a public alternative to the tools and products offered by industry giants. He stressed that scikit-learn was not the fruit of commercial interest and that decision-making about the project must always “come from the community.” The community ethos was underlined by other maintainers. Julien Jerphanion explained that only a minority of the scikit-learn community are paid for their contributions. “Even some of the maintainers are working on it in their spare time. The major part of the labour is based on benevolence of people working in their free time and not asking to get paid.” Community members perform many non-technical tasks, such as creating onboarding resources, helping the triaging team, and organising events. Many also make financial contributions in the form of micro-donations via NUMFOCUS. While “the major stuff is not funded through NUMFOCUS”, one maintainer explained that it has been useful in funding marketing, events, and so far one internship for underrepresented groups.
Beyond scikit-learn itself, several maintainers explained they care deeply about contributing upstream to the wider scientific Python ecosystem. During an on-site visit in September 2022, the maintainers had just returned from EuroSciPy, an annual gathering of European stakeholders in the scientific Python community. One maintainer spoke fondly of “putting faces to GitHub IDs.” Several maintainers mentioned that they felt a responsibility to support their dependencies, such as numpy and scipy, and that they collaborate with these projects in numerous ways beyond the software itself, including by applying for grants together. In addition, some of the maintainers had initiated “satellite projects,” such as joblib and skrub. However, their ability to support the ecosystem is constrained by their limited resources and hefty workload. Julien explained that while projects like scikit-learn, numpy, and scipy are very important, “There are many which are not visible to the public, for example OpenBLAS which is fundamental and only maintained by two people, who have very specific expertise. This is tremendous software work, but it is invisible to the public. Projects like this are cornerstones for many projects. They are the foundations; they are, like, infrastructure — you do not see them if there is no problem in your daily life.”
The maintainers also have personal interests in maintaining scikit-learn. Several maintainers explained they became maintainers because of their scientific interest in ML, and they like the scientific culture of the project and at Inria. Many of the maintainers have academic backgrounds, some work in research labs, and some teach at universities. One maintainer explained that while he may not become a millionaire contributing to open source, he had a comfortable life and enjoyed the challenge of the job, whilst another maintainer said that he enjoyed working at the intersection of statistical theory and software engineering, which was lacking in his previous job in industry. For many, being a maintainer of scikit-learn is also a source of pride, given the project’s reputation and impact in the field of ML/AI.
The diversity of incentives in the community is understood by the maintainers as a strength of the project, where decision-making is made based on the consensus of a diverse community. Adrin explained, “Even as a core developer, I don’t have enough credit to just be like, ‘Hey, we should do this.’ It’s a slow project with hundreds of 1000s of users. You can’t just change things.” Furthermore, the active involvement of the community was stressed as a key factor for the project’s sustainability. For example, Julien explained, “It is important to have people get paid for the work they do for scikit-learn. But if there’s only people that are getting paid, then something’s missing. I think what’s important it’s that there’s an understanding that an open source project is a commons and people can get involved if they want to.”
@@
4.2 From Inria to France’s AI strategy: The role of public funding in scikit-learn
The public sector has funded the project since its very beginning. While it initially started as scikits.learn during a Google Summer of Code project in 2007, it was abandoned and later exhumed by research fellows at Inria in 2010. To date, Inria has supported scikit-learn by employing staff, providing office space, and sponsoring events. A rough estimate of their sponsorship since the beginning is €1.5 million. “So, it was initially with public money from Inria from French public research grants that scikit-learn was developed,” explained François. Olivier Grisel, a maintainer, similarly explained that the public sector has played a long role in funding scikit-learn through Inria: “This is not new with the AI strategy.” Inria’s support role is not limited to funding. Through the Inria Foundation, which was established in 2017 to mobilise financial resources to support ambitious research projects, it has provided the institutional infrastructure for the scikit-learn consortium and its commercial sponsorship since 2018. One maintainer explained that while Inria may not be widely known in the community, its historical role in supporting scikit-learn cannot be overlooked.
The French government’s announcement that it would invest €32 million in the expansion of scikit-learn marked a new phase in the public sector’s involvement in the project. The grant is part of the strategy’s aim to cement France’s national competitiveness in AI and to increase AI adoption in the French economy. The strategy highlights that scikit-learn is an asset, which is used by millions of data scientists on a daily basis and which has been supported by Inria since its beginning. A government official explained that the funding for scikit-learn and related open source tools was intended to increase the adoption of AI throughout the country. By virtue of being open source, it was acknowledged that researchers and businesses beyond France would benefit, too. In addition to economic goals, the government official explained that digital sovereignty was a key consideration for funding scikit-learn and the digital commons more widely. This was underscored by secondary documents that interviewees shared with me that contextualised the government’s funding for scikit-learn as part of a broader strategy to develop sovereign digital commons in France and Europe. The same government official explained that there is a “real challenge” with the dependence on technologies provided by US Big Tech companies, even including open source frameworks like PyTorch and TensorFlow, in the field of AI and by investing in the digital commons the government aspired to address such dependencies.
@@ 4.2 From Inria to France’s AI strategy: The role of public funding in scikit-learn
Another government official explained the funding decision was informed by a multi-stakeholder process involving stakeholders from research and industry. “It was not just the government. It was a dialogue between several actors—with the AI strategy team, the Confiance IA project, [^1] Confiance IA is a government-sponsored programme to develop and industrialise trustworthy AI technologies. Inria, with scikit-learn, and so on.” However, the coordinators of the AI strategy faced hurdles within the government to get support for the grant: “Financing open source means financing a product that cannot be valued on the market and a priori it is not clear what its success will be. So, it’s very problematic in the sense that if the government finances it in a substantial way—here we are talking about €32 million—the government wants to know where the money is going and to achieve a measurable return on investment.” It took months of coordination and negotiation to eventually get the sign-off from the finance ministry. A maintainer explained that the sum would in fact be double of the announced sum, given a requirement for the scikit-learn team to double the sum, with one half from the private sector and the other half from other public bodies in Europe.
Several maintainers mentioned that they were pleased to receive support from the government. Julien commented, “scikit-learn is a global project, but it started at Inria, so it’s nice to get some recognition from the government. It’s a good signal that the French state is getting involved in funding projects like this.” François commended the government’s requirement to double the funding with equal halves stemming from public and private sources across the EU. “I think this is a good vision because we are a community with different interests. I think it would be dangerous for us to be exclusively funded by the private sector or to be exclusively funded by the French government, because we have many good contributors who are not French, and they could have the feeling that, ‘Okay, now it’s a French government project, they are buying it.”’ Meanwhile, Adrin highlighted that the funding had secured the sustainability of the project for at least five years, which meant that they could make long-term plans regarding recruitment and project development.
However, the government’s funding of scikit-learn has not been without its challenges. Some of the maintainers commented that the government’s focus on high-level public policy goals meant that some of their demands were removed from their daily practices or expertise as OSS developers. Initially, the government officials had neglected maintenance and requested that the scikit-learn team expand the library to deep learning (DL) to provide a competitive alternative to Meta’s PyTorch [^2] N.B. Meta donated PyTorch to the Linux Foundation in September 2022. and Google’s TensorFlow. The scikit-learn team pushed back against this request, reasoning that they should focus on their strength in ML and consolidate their maintenance capacity. The government was receptive to this feedback. A government official told me, “We acknowledge that scikit-learn is specialised in ML, and it should not compete against TensorFlow and PyTorch in DL.” Moreover, one maintainer mentioned that the operationalisation of government objectives “is not easy to do, and sometimes it really does not make sense.” For example, on the one hand, the government wants the scikit-learn team to develop open source data science tools, while on the other hand it requires that scikit-learn does not compete with French companies. In practice, this is not that easy to do because the tools that they will build will be open source alternatives to proprietary products and services offered by French companies. It was a tricky demand to balance these two expectations. Another challenge for the maintainers was the pace of delivery. One maintainer expressed that even 17 months after the initial announcement, they still had not received the first instalment of the grant. While scikit-learn can support itself through its mixed funding model, he explained that the slow pace of funding could be problematic for other projects in more precarious financial situations.
@@
While the interviewed sponsors saw the government’s funding positively, two representatives shared their concerns about the potential politicisation of the project owing to the French government’s substantial support for the project. One representative underlined the need to balance the French government’s concern for digital sovereignty with the norms of the global community, while another argued that imposing policy goals on the OSS project would not be received well by the community. These concerns were acknowledged by maintainers who I discussed these concerns with, who insisted that the community ethos of the project remained at the heart of their discussions with the government.
4.3 The muscle of industry: The role of the scikit-learn consortium
The scikit-learn project benefits from direct and indirect funding from companies that use scikit-learn in their business products and services. Some companies, such as Nvidia and Quansight Labs, have sponsored developers to contribute to scikit-learn. However, the most substantial funding has come from the scikit-learn consortium, which operates on an annual fee-based membership. François explained that its creation in 2018 addressed two pain points that the project faced: first, it would provide more stable funding for the project beyond individual research grants; and second, it would enable the project to offer permanent contracts to the maintainers, which was not possible previously at Inria. He explained the consortium offers companies “a way of giving back to the community and behaving responsibly towards a software tool that is part of their core business or their products.” In addition to financial support, the through the consortium’s Advisory Committee and the Technical Committee, companies advise on the project’s technical roadmap and share their industry use cases with the maintainers.
Companies pursue a combination of goals by joining the scikit-learn consortium. In both public blogs and interviews, companies cited reciprocity as a key motivation. Leo Dreyfus-Schmidt, the Dataiku representative, told me, “From our perspective, we think it’s a great opportunity to give back to the community because we are very much aware that we rely on open tools, especially scikit-learn.” A representative from another company added, “Companies pay millions in commercial licences, so why not support the OSS projects we use?” Companies also seek to capitalise on the reputational benefits of supporting scikit-learn, which in turn helps with hiring developers, who want to work at a company where OSS is considered important. François explained, “A lot of people come from this open source background and if you see that a company is promoting such an open source project, you might say, ‘Okay, maybe, I would be happy working with them. They could be okay with me contributing to OSS at work.”’ This incentive was viewd by some maintainers as more important for companies that did not already have a strong reputation in ML, since it would make them more credible and attractive to jobseekers.
Since all the companies in the consortium are users of scikit-learn, a key incentive for funding the project is to sustain its maintenance since it is a key dependency for their business. Mike McCarty from Nvidia shared that his concern about unmaintained dependencies “keeps me up at night” and cited NumPy as an example of a critical library that had been under-maintained for some time. In addition to maintenance, joining the consortium opens the door for a direct exchange between companies and the scikit-learn maintainers about industry use cases. Julien explained that he appreciates this exchange because “it shows culturally that those companies are not just using the software because it’s free, and it works; they are giving money to make the project sustainable.” Similarly, François commented, “[their] feedback is really rooted in something real. They have needs, they are working with customers on real projects, they need it to work.” Leo Dreyfus-Schmidt from Dataiku suggested, “I think that it’s a win-win for everybody here because putting myself in scikit-learn’s shoes, they don’t only get money, but they also get the perspective of people in industry, which for scikit-learn is a very different type of feedback from the one they get from the open source community. And it’s a win for all the consortium partners because it acts as a way to bring people together, which otherwise I wouldn’t know how that would happen.” Similarly, another representative explained, “It is good for the scikit-learn team to hear feedback from industry; not just small issues but how their tools are being used and what real world challenges are.” Several functionalities have been built into the scikit-learn library as a result of these exchanges.

### v4

@@ 3.2 Case presentation
scikit-learn is a Python library that implements ML algorithms for classification, regression, and clustering, as well as tools for data pre-processing, model evaluations, and data visualisation. Initially it was started as scikits.learn by David Courapeau during a Google Summer of Code project in 2007. After a dormant period, it was relaunched as scikit-learn by Fabian Pedregosa, Gaël Varoquaux, Alexandre Gramfort, and Vincent Michel with the backing of INRIA, with the first public release in February 2010. It has since become one of the most popular OSS libraries for ML [INRIA_2019_2020], described as “the Swiss army knife of ML” due to its widespread use [INRIA_2019_2020]. For example, in November 2023, it had 2,752 contributors and 61,085 GitHub stars [79]. The success and sustainability of this community-led OSS project is noteworthy, considering the extent of industry dominance in ML/AI research [6] and OSS ecosystems [80, 81].
scikit-learn’s sustainability is partly credited to its funding model, which has combined research grants, private sponsorships via the scikit-learn consortium, and community donations [82, 83]. The consortium was established under the Inria Foundation in 2018 in order to secure employment for maintainers, to sustain the quality of software, and to develop new features [83]. It comprises companies that pay annual membership fees. Since its founding, it has included Dataiku, Microsoft, Nvidia, Intel, AXA, Boston Consulting Group, BNP Paribas Cardif, Hugging Face, Nvidia, Fujitsu, and Chanel. Other companies, such as Quansight Labs, have indirectly funded the project by sponsoring maintainers. The project also benefits from donations from community members via NUMFOCUS and student projects have been sponsored by the Google Summer of Code programme. In November 2021, the French AI strategy, la “Stratégie IA”, announced a €32 million grant to support scikit-learn’s development [7]. This significant public investment alongside scikit-learn’s existing funding sources makes it a suitable case for the research aims of this exploratory case study.
@@ 3.3.2 Qualitative data analysis
Throughout the 17 months of data collection, an iterative, integrated approach was employed to identify common themes in the interview data, field notes, and secondary documents [85]. Given the exploratory aims of this study, the integrated approach combined the application of grounded theory analysis [86] and a deductive, organising framework informed by prior literature about the different interests and involvement models of individuals, companies, and governments in the development of OSS [37, 48, 72, 16]. Thematic codes were continuously developed as data was collected, and in turn translated into themes until the point of saturation [87]. The resulting themes were member-checked to ensure their accuracy and resonance with research participants [88]. Throughout the research process, a social identity map was maintained by the author to exercise reflexivity regarding the potential influence of personal biases stemming from the author’s views and relationships built with the scikit-learn team during the 17 months of data collection [89].
@@ 4 Results
The results are organised into three parts, sequentially responding to RQ1 on how the interests of scikit-learn have aligned with or differed from those of the maintainers, and RQ2 on the relative benefits and drawbacks of public and private funding from the perspective of the scikit-learn maintainers. Finally, practical lessons and recommendations for practitioners based on the funding model of scikit-learn RQ3 are discussed in the subsequent Discussion section.
4.1 Stewards of the Community: The Roles & Interests of the scikit-learn Maintainers
@@
scikit-learn’s maintainers are a diverse group employed by INRIA, public universities, and software companies. While some members of the core team work on scikit-learn full-time, others balance their roles as maintainers with their full-time jobs. Several maintainers also held academic positions at universities and research labs. The maintainers explained they perform a range of technical and social tasks, including reviewing code, mentoring contributors, and community building. François Goupil, the community manager, explained, “Of course, software development itself is the major part of the workload—going through issues, pull requests, and responding to community feedback—but lots of other work goes into the project. Community animation is a huge part. After all, it’s a community project”. In addition, sponsored developers from companies make focused contributions, such as the development of new features, which are agreed upon by the maintainers and the sponsoring companies. For example, at the time of this research, Nvidia sponsord a full-time maintainer who was building a plug-in to enable alternative computational backends for scikit-learn to enable compatibility with Nvidia’s GPU AI hardware accelerators.
Several maintainers expressed a strong sense of stewardship for the project, whose strength they viewed as rooted in its diverse community, comprising contributors from across the world (see Table 1) and companies (see Table 2). Julien Jerphanion, a maintainer, highlighted this community ethos, noting that, “the major part of the labour [in the project] is based on benevolence of people working in their free time and not asking to get paid”. Furthermore, Gaël Varoquaux, a co-founder and maintainer, explained that the community was at the heart of the project’s DNA and that major decisions about the future of the project remained community-driven despite their diverse funders. Similarly, Adrin Jalali, a maintainer, commented, “Even as a [maintainer], I don’t have enough credit to just be like, ‘Hey, we should do this.’ It’s a slow project with hundreds of 1000s of users. You can’t just change things”.
@@
In addition, the maintainers expressed their personal interests in maintaining scikit-learn. Several maintainers explained they became maintainers because of their scientific interest in ML, and they liked the scientific culture of the project. Many of the maintainers had academic backgrounds, some still worked in research labs, and some still taught at universities. For example, Julien Jerphanion said that he enjoyed working at the intersection of statistical theory and software engineering, which was lacking in his previous job in industry. In addition, an anonymous maintainer explained that while he may not become a millionaire, he had a comfortable life and enjoyed the challenge of the job. For many, being a maintainer of scikit-learn was also a source of pride, given the project’s reputation.
Beyond maintaining scikit-learn, several maintainers mentioned the importance for them to contribute to the wider scientific Python ecosystem and to support their dependencies, such as numpy and scipy. During my on-site visit in September 2022, the maintainers had just returned from EuroSciPy, an annual gathering of the European scientific Python community. One anonymous maintainer spoke fondly of “putting faces to GitHub IDs”. François Goupil explained that beyond collaborating on the code base, they collaborate with other projects in other ways, such as by applying for grants together. In addition, some of the maintainers had initiated “satellite projects”, such as joblib and skrub. However, their ability to support the ecosystem is constrained by their limited capacity. Julien Jerphanion explained that while projects like scikit-learn, numpy, and scipy are very important in ML/AI, “There are many which are not visible to the public, for example OpenBLAS which is fundamental and only maintained by two people, who have very specific expertise. This is tremendous software work, but it is invisible to the public. Projects like this are cornerstones for many projects. They are the foundations; they are, like, infrastructure — you do not see them if there is no problem in your daily life”.
4.2 Public Funding: From Research Grants at INRIA to “la Stratégie IA”
Public funding has been a cornerstone of scikit-learn’s development since its beginning. Starting as a Google Summer of Code project in 2007, the library was later revived by researchers at INRIA. Over the past 13 years, INRIA had provided an estimated €1.5 million in support through staff salaries, research grants, office space, computing resources, and event sponsorship. As Olivier Grisel, a maintainer, noted, “[Public funding] is not new with the AI strategy”. However, the announcement of a €32 million grant in France’s AI strategy, la “Stratégie IA”, supporting the expansion of scikit-learn and OSS tools for data science, marked a significant expansion of public funding for scikit-learn. A government official explained that the grant seeks to facilitate AI adoption across the economy and to enhance French competitiveness in AI R&D, while also supporting the “digital sovereignty” of France and Europe.
The maintainers were delighted to receive the funding and recognition from the government. Adrin Jalali highlighted the stability it provided for the project, enabling longer-term planning around hiring and the technical roadmap. Julien Jerphanion characterised it as “a good signal that the French state is getting involved in funding projects like this”. François Goupil particularly commended the requirement to seek matching funds from both public and private sources across the European Union, viewing this as an important safeguard for the project’s independence: “I think it would be dangerous for us to be exclusively funded by the private sector or to be exclusively funded by the French government, because we have many good contributors who are not French”.
At the same time, challenges have emerged in aligning broad policy goals with the day-to-day realities and expertise of the maintainers. Initial requests from officials to expand scikit-learn into deep learning to provide a “sovereign” alternative to PyTorch and TensorFlow were resisted, with maintainers insisting on focusing on the project’s core strength in classical ML and the need to allocate sufficient funding and capacity to maintenance. The government officials ultimately deferred to their expertise, with one official acknowledging, “We understand that scikit-learn is specialised in ML, and it should not compete against TensorFlow and PyTorch”.
However, other tensions have proven thornier. Requirements to simultaneously develop OSS tools for data science whilst not competing with domestic companies have left maintainers in a bind, as the resulting tools inherently provide open source alternatives to the proprietary offerings of companies. Maintainers have also expressed frustration with the slow pace of disbursement, with funds still not arriving 17 months after the announcement by the time of the last interview. While scikit-learn’s diverse funding model provided a buffer, they worried that such sluggishness could threaten the sustainability of more financially precarious OSS projects.
@@ 4.2 Public Funding: From Research Grants at INRIA to “la Stratégie IA”
Another official explained that the funding decision was informed by a multi-stakeholder process involving experts from research and industry. “It was not just the government. It was a dialogue between several actors—with the AI strategy team, the ‘Confiance IA’ project,111“Confiance IA” is a government-sponsored programme that aims to develop and industrialise trustworthy AI technologies in France. INRIA, with scikit-learn, and so on”. He confessed they had faced hurdles within the government: “Financing open source means financing a product that cannot be valued on the market and a priori it is not clear what its success will be. So, if [we want] the government to finance it in a substantial way—here we are talking about €32 million—the government wants to know where the money is going and to achieve a measurable return on investment”. Ultimately, it took months of coordination and negotiation to get the grant approved by the finance ministry.
@@
Two consortium members raised concerns that such a substantial government grant could politicise scikit-learn, potentially influencing it to prioritise French policy objectives over the interests of the community. One anonymous representative underlined the need to balance the French government’s concern about digital sovereignty with the norms of the global community, while another argued that imposing French policy priorities on the project would not be received well by the community. These concerns were acknowledged by the maintainers, who insisted that the community ethos of the project remained at the heart of their discussions with the government.
4.3 Private Sponsorship: The Role of the scikit-learn Consortium
scikit-learn has received significant support from the private sector through the scikit-learn consortium and sponsorship of maintainers by companies like Nvidia, HuggingFace, and QuantSight. By creating the consortium, François Goupil explained they had sought to provide a more stable source of funding beyond research grants and to hire full-time maintainers, which was not possible previously.
Companies in the consortium have pursued a range of social, economic, and technological goals. By joining the consortium, companies are able to support a software tool that they use, whilst building goodwill and a reputation as OSS advocates, which an important signal for hiring in a competitive labour market. François Goupil explained, “A lot of people come from this open source background and if you see that a company is promoting such an open source project, you might say, ’Okay, maybe, I would be happy working with them. They could be okay with me contributing to OSS at work”. He viewed this incentive as more important for companies that did not already have a strong reputation in ML/AI field, since it would make them more credible and attractive to job-seekers.
For their part, consortium members cited a mix of pragmatic and principled reasons for funding scikit-learn. On a basic level, it allows them to support the maintenance of a tool they use in their products and services. Mike McCarty from Nvidia shared that his concern about unmaintained dependencies “keeps [him] up at night”, citing numpy as a critical library that had been under-maintained for some time. Leo Dreyfus-Schmidt from Dataiku expressed a similar sentiment, who stated, “We are very much aware that we rely on open tools like scikit-learn” and highlighted the importance of “giv[ing] back to the community”. An anonymous consortium member was even more direct: “Companies pay millions in commercial licenses, so why not support the OSS projects we use¿‘
However, discussions with maintainers and consortium members underlined that consortium membership is not simply a charitable contribution. In exchange for their sponsorship, companies gain a voice in scikit-learn’s development through bodies like the Technical Committee and Advisory Board, which are venues for surfacing industry use cases and needs. Leo Dreyfus-Schmidt characterised this as a “win-win”, allowing scikit-learn to benefit from “the perspective of people in industry, which…is a very different type of feedback from the one they get from the open source community”, whilst giving commercial members direct input into the project.